# EMMA-500: Enhancing Massively Multilingual Adaptation of Large Language Models

### v1

@@ D.1 Commonsense Reasoning
Llama 2 7B 0.5667 0.4860 0.0224 0.5060 0.0224 0.6240 0.0217 0.6580 0.0212 0.5160 0.0224 0.5220 0.0224 0.5340 0.0223 0.5620 0.0222 0.5480 0.0223 0.6280 0.0216 0.6500 0.0214
Llama 2 7B Chat 0.5585 0.4780 0.0224 0.5080 0.0224 0.6240 0.0217 0.6720 0.0210 0.5060 0.0224 0.5220 0.0224 0.5060 0.0224 0.5500 0.0223 0.5520 0.0223 0.6120 0.0218 0.6140 0.0218
CodeLlama 2 7B 0.5469 0.4680 0.0223 0.5180 0.0224 0.5740 0.0221 0.6300 0.0216 0.5160 0.0224 0.4880 0.0224 0.5500 0.0223 0.5540 0.0223 0.5380 0.0223 0.5580 0.0222 0.6220 0.0217
LLaMAX Llama 2 7B 0.5438 0.4920 0.0224 0.5260 0.0224 0.5380 0.0223 0.5260 0.0224 0.5140 0.0224 0.5400 0.0223 0.5800 0.0221 0.5720 0.0221 0.5300 0.0223 0.5300 0.0223 0.6340 0.0216
LLaMAX Llama 2 7B Alpaca 0.5660 0.5120 0.0224 0.5420 0.0223 0.5720 0.0221 0.6100 0.0218 0.5240 0.0224 0.5500 0.0223 0.5700 0.0222 0.5640 0.0222 0.5520 0.0223 0.5520 0.0223 0.6780 0.0209
MaLA-500 Llama 2 10B v1 0.5309 0.4860 0.0224 0.5340 0.0223 0.5300 0.0223 0.5940 0.0220 0.5020 0.0224 0.5280 0.0223 0.5760 0.0221 0.5420 0.0223 0.5160 0.0224 0.5240 0.0224 0.5080 0.0224
MaLA-500 Llama 2 10B v2 0.5309 0.4860 0.0224 0.5340 0.0223 0.5300 0.0223 0.5940 0.0220 0.5020 0.0224 0.5280 0.0223 0.5760 0.0221 0.5420 0.0223 0.5160 0.0224 0.5240 0.0224 0.5080 0.0224
YaYi Llama 2 7B 0.5671 0.4880 0.0224 0.5080 0.0224 0.6260 0.0217 0.6700 0.0210 0.5060 0.0224 0.5320 0.0223 0.5520 0.0223 0.5420 0.0223 0.5540 0.0223 0.6320 0.0216 0.6280 0.0216
TowerBase Llama 2 7B 0.5633 0.4600 0.0223 0.5020 0.0224 0.6020 0.0219 0.7080 0.0204 0.5220 0.0224 0.5060 0.0224 0.5440 0.0223 0.5600 0.0222 0.5380 0.0223 0.5920 0.0220 0.6620 0.0212
TowerInstruct Llama 2 7B 0.5705 0.4880 0.0224 0.5160 0.0224 0.6200 0.0217 0.7100 0.0203 0.5200 0.0224 0.5100 0.0224 0.5420 0.0223 0.5640 0.0222 0.5460 0.0223 0.5860 0.0220 0.6740 0.0210
Occiglot Mistral 7B v0.1 0.5667 0.4720 0.0223 0.5140 0.0224 0.5700 0.0222 0.7460 0.0195 0.5200 0.0224 0.5160 0.0224 0.5720 0.0221 0.5580 0.0222 0.5440 0.0223 0.5520 0.0223 0.6700 0.0210
Occiglot Mistral 7B v0.1 Instruct 0.5655 0.4680 0.0223 0.5100 0.0224 0.5840 0.0221 0.7380 0.0197 0.5280 0.0223 0.5000 0.0224 0.5660 0.0222 0.5500 0.0223 0.5620 0.0222 0.5580 0.0222 0.6560 0.0213
BLOOM 7B 0.5689 0.4820 0.0224 0.5080 0.0224 0.6980 0.0206 0.5280 0.0223 0.5080 0.0224 0.5180 0.0224 0.5920 0.0220 0.5540 0.0223 0.5100 0.0224 0.7080 0.0204 0.6520 0.0213
BLOOMZ 7B 0.5487 0.4920 0.0224 0.5400 0.0223 0.6060 0.0219 0.5140 0.0224 0.5060 0.0224 0.5340 0.0223 0.5740 0.0221 0.5300 0.0223 0.5220 0.0224 0.5980 0.0219 0.6200 0.0217
mGPT 0.5504 0.5300 0.0223 0.4980 0.0224 0.5880 0.0220 0.5820 0.0221 0.5060 0.0224 0.5640 0.0222 0.5320 0.0223 0.5520 0.0223 0.5600 0.0222 0.6020 0.0219 0.5400 0.0223
mGPT 13B 0.5618 0.5080 0.0224 0.5180 0.0224 0.6260 0.0217 0.6080 0.0219 0.4820 0.0224 0.5800 0.0221 0.5480 0.0223 0.5280 0.0223 0.5680 0.0222 0.6340 0.0216 0.5800 0.0221
XGLM 7.5B 0.6064 0.6140 0.0218 0.5740 0.0221 0.6940 0.0206 0.6360 0.0215 0.4880 0.0224 0.6000 0.0219 0.5460 0.0223 0.5940 0.0220 0.5840 0.0221 0.7020 0.0205 0.6380 0.0215
YaYi 7B 0.5664 0.5040 0.0224 0.5300 0.0223 0.6340 0.0216 0.5180 0.0224 0.5180 0.0224 0.5540 0.0223 0.5620 0.0222 0.5460 0.0223 0.5200 0.0224 0.6640 0.0211 0.6800 0.0209
Llama 3 8B 0.6171 0.5340 0.0223 0.5280 0.0223 0.7140 0.0202 0.7160 0.0202 0.5120 0.0224 0.5800 0.0221 0.6000 0.0219 0.5860 0.0220 0.6240 0.0217 0.7160 0.0202 0.6780 0.0209
Llama 3.1 8B 0.6171 0.5300 0.0223 0.5380 0.0223 0.7160 0.0202 0.7260 0.0200 0.4880 0.0224 0.5540 0.0223 0.6120 0.0218 0.5780 0.0221 0.6180 0.0218 0.7240 0.0200 0.7040 0.0204
Gemma 7B 0.6364 0.5920 0.0220 0.5480 0.0223 0.7200 0.0201 0.7280 0.0199 0.5000 0.0224 0.6060 0.0219 0.6160 0.0218 0.6040 0.0219 0.6560 0.0213 0.7420 0.0196 0.6880 0.0207
Gemma 2 9B 0.6633 0.6380 0.0215 0.5280 0.0223 0.7780 0.0186 0.7580 0.0192 0.5040 0.0224 0.6380 0.0215 0.6360 0.0215 0.6380 0.0215 0.6740 0.0210 0.7660 0.0190 0.7380 0.0197
Qwen 2 7B 0.6031 0.5080 0.0224 0.5080 0.0224 0.7060 0.0204 0.7140 0.0202 0.5040 0.0224 0.5240 0.0224 0.5280 0.0223 0.6100 0.0218 0.5720 0.0221 0.6940 0.0206 0.7660 0.0190
Qwen 1.5 7B 0.5944 0.5200 0.0224 0.5300 0.0223 0.6440 0.0214 0.6520 0.0213 0.5100 0.0224 0.5260 0.0224 0.5580 0.0222 0.5760 0.0221 0.5880 0.0220 0.6920 0.0207 0.7420 0.0196
EMMA-500 Llama 2 7B 0.6311 0.6140 0.0218 0.5800 0.0221 0.7420 0.0196 0.6940 0.0206 0.5240 0.0224 0.6620 0.0212 0.6000 0.0219 0.5560 0.0222 0.6200 0.0217 0.7020 0.0205 0.6480 0.0214
Table 22: 0-shot results (ACC) on XStoryCloze in all languages
@@
Llama 2 7B 0.5755 0.4990 0.0129 0.7704 0.0108 0.6737 0.0121 0.5036 0.0129 0.5374 0.0128 0.5923 0.0126 0.4805 0.0129 0.6300 0.0124 0.5050 0.0129 0.5433 0.0128 0.5956 0.0126
Llama 2 7B Chat 0.5841 0.5050 0.0129 0.7869 0.0105 0.6711 0.0121 0.5083 0.0129 0.5407 0.0128 0.5963 0.0126 0.4864 0.0129 0.6552 0.0122 0.5202 0.0129 0.5334 0.0128 0.6221 0.0125
CodeLlama 2 7B 0.5568 0.5010 0.0129 0.7148 0.0116 0.6340 0.0124 0.5043 0.0129 0.4970 0.0129 0.5586 0.0128 0.4937 0.0129 0.5923 0.0126 0.5003 0.0129 0.5374 0.0128 0.5917 0.0126
LLaMAX Llama 2 7B 0.6036 0.5890 0.0127 0.7551 0.0111 0.6525 0.0123 0.5447 0.0128 0.5817 0.0127 0.6062 0.0126 0.5248 0.0129 0.6122 0.0125 0.5718 0.0127 0.5930 0.0126 0.6082 0.0126
LLaMAX Llama 2 7B Alpaca 0.6383 0.6036 0.0126 0.8147 0.0100 0.7068 0.0117 0.5486 0.0128 0.6214 0.0125 0.6645 0.0122 0.5381 0.0128 0.6744 0.0121 0.6016 0.0126 0.5930 0.0126 0.6545 0.0122
MaLA-500 Llama 2 10B v1 0.5307 0.4818 0.0129 0.7353 0.0114 0.6241 0.0125 0.4990 0.0129 0.4765 0.0129 0.4792 0.0129 0.4626 0.0128 0.5493 0.0128 0.4871 0.0129 0.5261 0.0128 0.5169 0.0129
MaLA-500 Llama 2 10B v2 0.5307 0.4818 0.0129 0.7353 0.0114 0.6241 0.0125 0.4990 0.0129 0.4765 0.0129 0.4792 0.0129 0.4626 0.0128 0.5493 0.0128 0.4871 0.0129 0.5261 0.0128 0.5169 0.0129
YaYi Llama 2 7B 0.5842 0.4997 0.0129 0.7909 0.0105 0.6870 0.0119 0.5063 0.0129 0.5427 0.0128 0.6142 0.0125 0.4745 0.0129 0.6479 0.0123 0.5003 0.0129 0.5394 0.0128 0.6234 0.0125
TowerBase Llama 2 7B 0.5778 0.4917 0.0129 0.7723 0.0108 0.6982 0.0118 0.5076 0.0129 0.5288 0.0128 0.5831 0.0127 0.4838 0.0129 0.6704 0.0121 0.5036 0.0129 0.5314 0.0128 0.5850 0.0127
TowerInstruct Llama 2 7B 0.5924 0.4931 0.0129 0.8087 0.0101 0.7161 0.0116 0.5069 0.0129 0.5295 0.0128 0.5956 0.0126 0.4871 0.0129 0.6936 0.0119 0.5149 0.0129 0.5414 0.0128 0.6300 0.0124
Occiglot Mistral 7B v0.1 0.5810 0.5129 0.0129 0.7737 0.0108 0.7340 0.0114 0.5208 0.0129 0.5149 0.0129 0.5864 0.0127 0.4798 0.0129 0.6294 0.0124 0.4983 0.0129 0.5314 0.0128 0.6089 0.0126
Occiglot Mistral 7B v0.1 Instruct 0.5939 0.5268 0.0128 0.7942 0.0104 0.7419 0.0113 0.5301 0.0128 0.5275 0.0128 0.6036 0.0126 0.4825 0.0129 0.6506 0.0123 0.5043 0.0129 0.5381 0.0128 0.6334 0.0124
BLOOM 7B 0.5930 0.5857 0.0127 0.7055 0.0117 0.6618 0.0122 0.5725 0.0127 0.6042 0.0126 0.6453 0.0123 0.4891 0.0129 0.5275 0.0128 0.5394 0.0128 0.5731 0.0127 0.6188 0.0125
BLOOMZ 7B 0.5712 0.5652 0.0128 0.7300 0.0114 0.6459 0.0123 0.5109 0.0129 0.5764 0.0127 0.5533 0.0128 0.4825 0.0129 0.5215 0.0129 0.5215 0.0129 0.5817 0.0127 0.5943 0.0126
mGPT 0.5443 0.4931 0.0129 0.5996 0.0126 0.5546 0.0128 0.5460 0.0128 0.5275 0.0128 0.5314 0.0128 0.5122 0.0129 0.5665 0.0128 0.5500 0.0128 0.5725 0.0127 0.5341 0.0128
mGPT 13B 0.5644 0.5162 0.0129 0.6420 0.0123 0.5864 0.0127 0.5539 0.0128 0.5506 0.0128 0.5811 0.0127 0.5122 0.0129 0.5943 0.0126 0.5460 0.0128 0.5764 0.0127 0.5493 0.0128
XGLM 7.5B 0.6075 0.5612 0.0128 0.6982 0.0118 0.6386 0.0124 0.5771 0.0127 0.5884 0.0127 0.6300 0.0124 0.5705 0.0127 0.6340 0.0124 0.5936 0.0126 0.6023 0.0126 0.5890 0.0127
YaYi 7B 0.6067 0.6181 0.0125 0.7432 0.0112 0.6942 0.0119 0.5606 0.0128 0.6367 0.0124 0.6241 0.0125 0.4924 0.0129 0.5215 0.0129 0.5361 0.0128 0.5791 0.0127 0.6678 0.0121
Llama 3 8B 0.6341 0.5864 0.0127 0.7869 0.0105 0.7062 0.0117 0.5579 0.0128 0.6281 0.0124 0.6592 0.0122 0.5109 0.0129 0.6876 0.0119 0.5639 0.0128 0.6300 0.0124 0.6578 0.0122
Llama 3.1 8B 0.6358 0.5910 0.0127 0.7816 0.0106 0.7081 0.0117 0.5533 0.0128 0.6327 0.0124 0.6803 0.0120 0.5242 0.0129 0.6863 0.0119 0.5592 0.0128 0.6115 0.0125 0.6658 0.0121
Gemma 7B 0.6501 0.6042 0.0126 0.8015 0.0103 0.7062 0.0117 0.5758 0.0127 0.6492 0.0123 0.6764 0.0120 0.5228 0.0129 0.7055 0.0117 0.6214 0.0125 0.6327 0.0124 0.6559 0.0122
Gemma 2 9B 0.6767 0.6525 0.0123 0.8015 0.0103 0.7426 0.0113 0.6016 0.0126 0.6691 0.0121 0.7134 0.0116 0.5513 0.0128 0.7373 0.0113 0.6373 0.0124 0.6479 0.0123 0.6896 0.0119
Qwen 2 7B 0.6146 0.5996 0.0126 0.7895 0.0105 0.6976 0.0118 0.5202 0.0129 0.5797 0.0127 0.6439 0.0123 0.4897 0.0129 0.6956 0.0118 0.5142 0.0129 0.5407 0.0128 0.6903 0.0119
Qwen 1.5 7B 0.5985 0.5539 0.0128 0.7816 0.0106 0.6830 0.0120 0.5189 0.0129 0.5566 0.0128 0.6287 0.0124 0.4924 0.0129 0.6327 0.0124 0.5169 0.0129 0.5394 0.0128 0.6797 0.0120
EMMA-500 Llama 2 7B 0.6638 0.6625 0.0122 0.7644 0.0109 0.7002 0.0118 0.6473 0.0123 0.6492 0.0123 0.6863 0.0119 0.5791 0.0127 0.6850 0.0120 0.6473 0.0123 0.6466 0.0123 0.6340 0.0124
Table 23: 0-shot results (ACC) on XWinograd in all languages

### v2

@@ Appendix E Detailed Results
Llama 2 7B 56.67 48.60 2.24 50.60 2.24 62.40 2.17 65.80 2.12 51.60 2.24 52.20 2.24 53.40 2.23 56.20 2.22 54.80 2.23 62.80 2.16 65.00 2.14
Llama 2 7B Chat 55.85 47.80 2.24 50.80 2.24 62.40 2.17 67.20 2.10 50.60 2.24 52.20 2.24 50.60 2.24 55.00 2.23 55.20 2.23 61.20 2.18 61.40 2.18
CodeLlama 2 7B 54.69 46.80 2.23 51.80 2.24 57.40 2.21 63.00 2.16 51.60 2.24 48.80 2.24 55.00 2.23 55.40 2.23 53.80 2.23 55.80 2.22 62.20 2.17
LLaMAX Llama 2 7B 54.38 49.20 2.24 52.60 2.24 53.80 2.23 52.60 2.24 51.40 2.24 54.00 2.23 58.00 2.21 57.20 2.21 53.00 2.23 53.00 2.23 63.40 2.16
LLaMAX Llama 2 7B Alpaca 56.60 51.20 2.24 54.20 2.23 57.20 2.21 61.00 2.18 52.40 2.24 55.00 2.23 57.00 2.22 56.40 2.22 55.20 2.23 55.20 2.23 67.80 2.09
MaLA-500 Llama 2 10B v1 53.09 48.60 2.24 53.40 2.23 53.00 2.23 59.40 2.20 50.20 2.24 52.80 2.23 57.60 2.21 54.20 2.23 51.60 2.24 52.40 2.24 50.80 2.24
MaLA-500 Llama 2 10B v2 53.09 48.60 2.24 53.40 2.23 53.00 2.23 59.40 2.20 50.20 2.24 52.80 2.23 57.60 2.21 54.20 2.23 51.60 2.24 52.40 2.24 50.80 2.24
YaYi Llama 2 7B 56.71 48.80 2.24 50.80 2.24 62.60 2.17 67.00 2.10 50.60 2.24 53.20 2.23 55.20 2.23 54.20 2.23 55.40 2.23 63.20 2.16 62.80 2.16
TowerBase Llama 2 7B 56.33 46.00 2.23 50.20 2.24 60.20 2.19 70.80 2.04 52.20 2.24 50.60 2.24 54.40 2.23 56.00 2.22 53.80 2.23 59.20 2.20 66.20 2.12
TowerInstruct Llama 2 7B 57.05 48.80 2.24 51.60 2.24 62.00 2.17 71.00 2.03 52.00 2.24 51.00 2.24 54.20 2.23 56.40 2.22 54.60 2.23 58.60 2.20 67.40 2.10
Occiglot Mistral 7B v0.1 56.67 47.20 2.23 51.40 2.24 57.00 2.22 74.60 1.95 52.00 2.24 51.60 2.24 57.20 2.21 55.80 2.22 54.40 2.23 55.20 2.23 67.00 2.10
Occiglot Mistral 7B v0.1 Instruct 56.55 46.80 2.23 51.00 2.24 58.40 2.21 73.80 1.97 52.80 2.23 50.00 2.24 56.60 2.22 55.00 2.23 56.20 2.22 55.80 2.22 65.60 2.13
BLOOM 7B 56.89 48.20 2.24 50.80 2.24 69.80 2.06 52.80 2.23 50.80 2.24 51.80 2.24 59.20 2.20 55.40 2.23 51.00 2.24 70.80 2.04 65.20 2.13
BLOOMZ 7B 54.87 49.20 2.24 54.00 2.23 60.60 2.19 51.40 2.24 50.60 2.24 53.40 2.23 57.40 2.21 53.00 2.23 52.20 2.24 59.80 2.19 62.00 2.17
mGPT 55.04 53.00 2.23 49.80 2.24 58.80 2.20 58.20 2.21 50.60 2.24 56.40 2.22 53.20 2.23 55.20 2.23 56.00 2.22 60.20 2.19 54.00 2.23
mGPT 13B 56.18 50.80 2.24 51.80 2.24 62.60 2.17 60.80 2.19 48.20 2.24 58.00 2.21 54.80 2.23 52.80 2.23 56.80 2.22 63.40 2.16 58.00 2.21
YaYi 7B 56.64 50.40 2.24 53.00 2.23 63.40 2.16 51.80 2.24 51.80 2.24 55.40 2.23 56.20 2.22 54.60 2.23 52.00 2.24 66.40 2.11 68.00 2.09
Aya 23 8B 55.13 50.20 2.24 52.20 2.24 57.00 2.22 60.40 2.19 50.40 2.24 52.20 2.24 55.60 2.22 52.60 2.24 59.60 2.20 58.80 2.20 57.40 2.21
Aya expanse 8B 56.38 50.80 2.24 51.60 2.24 61.20 2.18 63.40 2.16 51.60 2.24 53.60 2.23 55.20 2.23 50.20 2.24 57.80 2.21 60.20 2.19 64.60 2.14
Llama 3 8B 61.71 53.40 2.23 52.80 2.23 71.40 2.02 71.60 2.02 51.20 2.24 58.00 2.21 60.00 2.19 58.60 2.20 62.40 2.17 71.60 2.02 67.80 2.09
Llama 3.1 8B 61.71 53.00 2.23 53.80 2.23 71.60 2.02 72.60 2.00 48.80 2.24 55.40 2.23 61.20 2.18 57.80 2.21 61.80 2.18 72.40 2.00 70.40 2.04
Gemma 7B 63.64 59.20 2.20 54.80 2.23 72.00 2.01 72.80 1.99 50.00 2.24 60.60 2.19 61.60 2.18 60.40 2.19 65.60 2.13 74.20 1.96 68.80 2.07
Gemma 2 9B 66.33 63.80 2.15 52.80 2.23 77.80 1.86 75.80 1.92 50.40 2.24 63.80 2.15 63.60 2.15 63.80 2.15 67.40 2.10 76.60 1.90 73.80 1.97
Qwen 2 7B 60.31 50.80 2.24 50.80 2.24 70.60 2.04 71.40 2.02 50.40 2.24 52.40 2.24 52.80 2.23 61.00 2.18 57.20 2.21 69.40 2.06 76.60 1.90
Qwen 1.5 7B 59.44 52.00 2.24 53.00 2.23 64.40 2.14 65.20 2.13 51.00 2.24 52.60 2.24 55.80 2.22 57.60 2.21 58.80 2.20 69.20 2.07 74.20 1.96
EMMA-500 Llama 2 7B 63.11 61.40 2.18 58.00 2.21 74.20 1.96 69.40 2.06 52.40 2.24 66.20 2.12 60.00 2.19 55.60 2.22 62.00 2.17 70.20 2.05 64.80 2.14
Table 19: 0-shot results (ACC %) on XStoryCloze in all languages
@@
Llama 2 7B 57.55 49.90 1.29 77.04 1.08 67.37 1.21 50.36 1.29 53.74 1.28 59.23 1.26 48.05 1.29 63.00 1.24 50.50 1.29 54.33 1.28 59.56 1.26
Llama 2 7B Chat 58.41 50.50 1.29 78.69 1.05 67.11 1.21 50.83 1.29 54.07 1.28 59.63 1.26 48.64 1.29 65.52 1.22 52.02 1.29 53.34 1.28 62.21 1.25
CodeLlama 2 7B 55.68 50.10 1.29 71.48 1.16 63.40 1.24 50.43 1.29 49.70 1.29 55.86 1.28 49.37 1.29 59.23 1.26 50.03 1.29 53.74 1.28 59.17 1.26
LLaMAX Llama 2 7B 60.36 58.90 1.27 75.51 1.11 65.25 1.23 54.47 1.28 58.17 1.27 60.62 1.26 52.48 1.29 61.22 1.25 57.18 1.27 59.30 1.26 60.82 1.26
LLaMAX Llama 2 7B Alpaca 63.83 60.36 1.26 81.47 1.00 70.68 1.17 54.86 1.28 62.14 1.25 66.45 1.22 53.81 1.28 67.44 1.21 60.16 1.26 59.30 1.26 65.45 1.22
MaLA-500 Llama 2 10B v1 53.07 48.18 1.29 73.53 1.14 62.41 1.25 49.90 1.29 47.65 1.29 47.92 1.29 46.26 1.28 54.93 1.28 48.71 1.29 52.61 1.28 51.69 1.29
MaLA-500 Llama 2 10B v2 53.07 48.18 1.29 73.53 1.14 62.41 1.25 49.90 1.29 47.65 1.29 47.92 1.29 46.26 1.28 54.93 1.28 48.71 1.29 52.61 1.28 51.69 1.29
YaYi Llama 2 7B 58.42 49.97 1.29 79.09 1.05 68.70 1.19 50.63 1.29 54.27 1.28 61.42 1.25 47.45 1.29 64.79 1.23 50.03 1.29 53.94 1.28 62.34 1.25
TowerBase Llama 2 7B 57.78 49.17 1.29 77.23 1.08 69.82 1.18 50.76 1.29 52.88 1.28 58.31 1.27 48.38 1.29 67.04 1.21 50.36 1.29 53.14 1.28 58.50 1.27
TowerInstruct Llama 2 7B 59.24 49.31 1.29 80.87 1.01 71.61 1.16 50.69 1.29 52.95 1.28 59.56 1.26 48.71 1.29 69.36 1.19 51.49 1.29 54.14 1.28 63.00 1.24
Occiglot Mistral 7B v0.1 58.10 51.29 1.29 77.37 1.08 73.40 1.14 52.08 1.29 51.49 1.29 58.64 1.27 47.98 1.29 62.94 1.24 49.83 1.29 53.14 1.28 60.89 1.26
Occiglot Mistral 7B v0.1 Instruct 59.39 52.68 1.28 79.42 1.04 74.19 1.13 53.01 1.28 52.75 1.28 60.36 1.26 48.25 1.29 65.06 1.23 50.43 1.29 53.81 1.28 63.34 1.24
BLOOM 7B 59.30 58.57 1.27 70.55 1.17 66.18 1.22 57.25 1.27 60.42 1.26 64.53 1.23 48.91 1.29 52.75 1.28 53.94 1.28 57.31 1.27 61.88 1.25
BLOOMZ 7B 57.12 56.52 1.28 73.00 1.14 64.59 1.23 51.09 1.29 57.64 1.27 55.33 1.28 48.25 1.29 52.15 1.29 52.15 1.29 58.17 1.27 59.43 1.26
mGPT 54.43 49.31 1.29 59.96 1.26 55.46 1.28 54.60 1.28 52.75 1.28 53.14 1.28 51.22 1.29 56.65 1.28 55.00 1.28 57.25 1.27 53.41 1.28
mGPT 13B 56.44 51.62 1.29 64.20 1.23 58.64 1.27 55.39 1.28 55.06 1.28 58.11 1.27 51.22 1.29 59.43 1.26 54.60 1.28 57.64 1.27 54.93 1.28
YaYi 7B 60.67 61.81 1.25 74.32 1.12 69.42 1.19 56.06 1.28 63.67 1.24 62.41 1.25 49.24 1.29 52.15 1.29 53.61 1.28 57.91 1.27 66.78 1.21
Aya 23 8B 60.93 62.61 1.25 74.59 1.12 67.31 1.21 50.96 1.29 64.26 1.23 66.64 1.21 47.72 1.29 68.70 1.19 50.36 1.29 54.00 1.28 63.14 1.24
Aya expanse 8B 64.80 69.42 1.19 80.41 1.02 74.26 1.13 51.36 1.29 68.50 1.20 72.47 1.15 48.25 1.29 73.86 1.13 51.62 1.29 55.13 1.28 67.57 1.20
Llama 3 8B 63.41 58.64 1.27 78.69 1.05 70.62 1.17 55.79 1.28 62.81 1.24 65.92 1.22 51.09 1.29 68.76 1.19 56.39 1.28 63.00 1.24 65.78 1.22
Llama 3.1 8B 63.58 59.10 1.27 78.16 1.06 70.81 1.17 55.33 1.28 63.27 1.24 68.03 1.20 52.42 1.29 68.63 1.19 55.92 1.28 61.15 1.25 66.58 1.21
Gemma 7B 65.01 60.42 1.26 80.15 1.03 70.62 1.17 57.58 1.27 64.92 1.23 67.64 1.20 52.28 1.29 70.55 1.17 62.14 1.25 63.27 1.24 65.59 1.22
Gemma 2 9B 67.67 65.25 1.23 80.15 1.03 74.26 1.13 60.16 1.26 66.91 1.21 71.34 1.16 55.13 1.28 73.73 1.13 63.73 1.24 64.79 1.23 68.96 1.19
Qwen 2 7B 61.46 59.96 1.26 78.95 1.05 69.76 1.18 52.02 1.29 57.97 1.27 64.39 1.23 48.97 1.29 69.56 1.18 51.42 1.29 54.07 1.28 69.03 1.19
Qwen 1.5 7B 59.85 55.39 1.28 78.16 1.06 68.30 1.20 51.89 1.29 55.66 1.28 62.87 1.24 49.24 1.29 63.27 1.24 51.69 1.29 53.94 1.28 67.97 1.20
EMMA-500 Llama 2 7B 66.38 66.25 1.22 76.44 1.09 70.02 1.18 64.73 1.23 64.92 1.23 68.63 1.19 57.91 1.27 68.50 1.20 64.73 1.23 64.66 1.23 63.40 1.24
Table 20: 0-shot results (ACC %) on XWinograd in all languages